OpenStudy (anonymous):

Help Solve each system by elimination -2-18y=2x -4x=-9y-23

OpenStudy (anonymous):

what r ur options?

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

do have any options in the answrs like (4,5) or something

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now what you need to do is that rearrange the two equations in such a way that the x and y terms r on one side

OpenStudy (anonymous):

that's what im stuck on it

OpenStudy (anonymous):

okay first number the equations for easy reference

OpenStudy (anonymous):

-2-18y=2x -4x=-9y-23 so for 1 -2x - 18y = +18 eqn 1 and for 2 14x + 9y = -23 eqn 2

OpenStudy (anonymous):

@glitterythings

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

get what

OpenStudy (anonymous):

the equations

OpenStudy (anonymous):

i have rearranged the equations

OpenStudy (anonymous):

I don't get how

OpenStudy (anonymous):

@glitterythings

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I just need to know how to rearrange them and to see what numbers can go into y

OpenStudy (anonymous):

can cross out y

OpenStudy (anonymous):

okay im ready to show you how to rearrange okay.

OpenStudy (anonymous):

@ChiefArnav shouldnt the equations be 2x + 18y = -2 and 4x - 9y = 23

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes ^^

OpenStudy (anonymous):

okay @Rohith15 thx for correcting my error

OpenStudy (anonymous):

okay @glitterythings -2-18y=2x so now get 2x on the right hand side so 2x -2 - 18y 2x- 18y = 2 got this much

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

now same for the second one do it and show will not do evrything for you

OpenStudy (anonymous):

okay okay

OpenStudy (anonymous):

im really confused

OpenStudy (anonymous):

okay for the second one -4x=-9y-23 moving y to the right hand side so 14x + 9y = -23

OpenStudy (anonymous):

okay I get it thanks!

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

@glitterythings its now 14x its -4x

OpenStudy (anonymous):

not*